# AniMer+: Unified Pose and Shape Estimation Across Mammalia and Aves via Family-Aware Transformer

### v1

@@ I Introduction
Accurate estimation of animal pose and shape from images is essential for capturing animal behavior, biomechanics, and interactions with environment. It provides vital insights for diverse fields including animal welfare, agriculture, ecology, and life sciences. Integrating geometric and appearance information from widely varying species, such as birds and mammals, into a unified deep neural network represents a crucial step towards comprehensively interpreting the intricate and dynamic animal world. Significant efforts in this domain involve estimating pose and shape parameters using articulated animal templates. For example, the well-known SMAL model [1] primarily targets mammals, while the AVES parametric model [2] focuses on birds. Although extensive research has been conducted on specific species or families, such as horses [3, 4, 5], zebras [6], dogs [7, 8, 9, 10, 11, 12] and birds [2], investigations into other mammal species, including cats, cows or hippos, remain relatively underexplored, and few previous methods try to predict the meshes for animals of different classes (e.g., mammalia and aves) in a single network.
@@
Reconstructing multiple species within a single network is hindered by two primary challenges: the limited capacity of backbones and the scarcity of multi-species datasets annotated with corresponding parametric model labels. Recent advancements in human mesh recovery via the SMPL model [13] demonstrate that combining high-capacity Transformer [14] backbones with large-scale datasets significantly enhances pose and shape estimation accuracy across diverse settings [15, 16, 17]. However, this effective paradigm has yet to be fully explored within animal studies. While  [18] introduces the large-scale Animal3D dataset featuring SMAL annotations, their work primarily utilizes CNN-based networks like HMR [19] and WLDO [11].
@@
Figure 1: Statistics and representative samples of the CtrlAVES3D and CtrlAni3D datasets. (a) and (b) represent samples from the CtrlAVES3D dataset and the CtrlAni3D dataset, respectively. For each image pair, the left side displays the generated animal image whose background comes from either COCO  [20] or AI-synthesis, and the right side presents the rendered mesh label. Note that the synthesis process naturally considers generating truncated images. (c) shows the statistics of both datasets. More details about these two datasets can be found in Table I, Table II and Sec. IV.
@@
To recap, our previous work, AniMer [21], proposes a systematic approach that pursues accurate animal pose and shape estimation by leveraging family-aware Transformer. The success of AniMer is built on two key scaling factors: a scaled backbone and a scaled dataset. Architecturally, AniMer integrates a high-capacity Vision Transformer (ViT) backbone with a Transformer-based decoder to regress SMAL parameters, a paradigm proven effective in human mesh recovery but previously untested for animal shape and pose estimation. To unify the discriminative understanding of diverse animal shapes within a single framework, AniMer introduces a novel animal family supervised contrastive learning scheme. It employs a learnable class token to explicitly enhance the model’s ability to distinguish between the unique shapes of different mammal families.
@@
A high-capacity model’s effectiveness is intrinsically linked to the quality and quantity of its training data. However, acquiring a large number of animal images with full 3D annotations is difficult, in which situations synthetic datasets can mitigate the scarcity of real data. The coarse texture quality and sophisticated lighting and shadow control hinder the usability of traditional Computer Graphics (CG) generated animal images. These limitations motivate our usage of ControlNet [22] to create high-quality AI hallucinated images conditioned on SMAL structures. Compared to CG-based, this principle effectively bridges the domain gap between parametric labels and images, and helps to create a novel, large-scale dataset named CtrlAni3D with minimal human labor involved. Specifically, we prompt ControlNet with textual descriptions of animal behaviors and rendered SMAL mask and depth images, resulting in highly realistic visual outputs, as illustrated in Fig. 1. To ensure dataset quality, we use SAM2.0 [23] together with manual verification to filter inappropriate images. Finally, CtrlAni3D comprises 9711 images annotated with pixel-aligned SMAL meshes, and its scalable nature allows for further expansion. We end up with scaled training datasets by aggregating most available open-sourced quadrupedal datasets and our CtrlAni3D dataset, resulting in a comprehensive set of 41.3k images annotated with either 3D meshes or 2D keypoints.
@@
Despite these advances, the SMAL model is insufficient for animals like birds, whose skeletal structures fundamentally differ from mammals’. Training a shared model naively on both mammals and birds risks “averaging out” crucial anatomical features and suffers from severe depth ambiguities in single-view reconstruction, exacerbated by the fact that existing avian datasets provide only 2D annotations. Inspired by the success of Mixture-of-Experts (MoE) architectures [24, 25, 26] in handling diverse tasks, we propose AniMer+, an extension of AniMer incorporating an MoE design to better unify different animal taxa. We adapt the ViT encoder by partitioning its fully-connected (FC) layers into taxa-specific layers (capturing expert quadruped/avian knowledge) and a taxa-shared layer (learning common semantic features). The taxa-specific layers process features from the corresponding taxa, while the taxa-shared layer receives features from all taxa. This architecture enables effective handling of both taxonomic groups within a single, unified framework. To enable robust 3D training for birds, we apply our generation pipeline to create CtrlAVES3D, the first large-scale avian dataset with 3D AVES model annotations, thereby mitigating the single-view ambiguity issue.
@@
To rigorously evaluate the efficacy of our AniMer and AniMer+ models, as well as our dataset generation pipeline, we conduct extensive empirical studies. For a clarification purpose, we name AniMer trained on mammals only, birds only and both mammals and birds respectively as AniMer-M, AniMer-A and AniMer-AM. Our findings indicate that AniMer, when trained on the complete multi-species mammal datasets (AniMer-M), significantly outperforms CNN-based methods such as HMR and WLDO [11] which serve as baselines in the Animal3D research. Similarly, AniMer-A largely surpasses previous methods on bird datasets. Meanwhile, through ablation studies, we verify that the synthesized datasets enhance the generalizability of AniMer on the out-of-domain (OOD) Animal Kingdom dataset [27] and the OOD Cow Bird dataset [28], which are unseen during training. In addition, ablation studies also prove that the animal family-supervised contrastive learning scheme improves pose and shape estimation precision. Finally, the joint training of AniMer+ on both mammals and birds outperforms AniMer-AM, demonstrating the effectiveness of the MoE design.
@@
An earlier version of this work is AniMer [21], which was published in CVPR 2025. In this paper, we have significantly enhanced AniMer in three aspects. First, we create the first large-scale bird dataset with 3D annotations, named CtrlAVES3D, which fulfills the lack of 3D annotations in existing avian datasets and mitigates depth ambiguity in the monocular bird reconstruction task. Second, we propose AniMer+ by incorporating an MoE design to better unify different animal taxa, which outperforms AniMer when trained simultaneously on mammals and birds. Third, we conduct comparative and ablation experiments on avian species, establishing state-of-the-art performance across multiple avian datasets. Based on these extensions and in conjunction with the dataset generation pipeline, AniMer+ has the potential to generalize to any animal groups (e.g., fishes, insects, rodents or primates) that can be represented by parametric models.
@@ II Related Works
Pose and Shape Estimation for Mammals and Birds. Compared to monocular 2D/3D animal keypoint estimation methods [29, 30, 31, 32, 33], monocular surface estimation methods can provide both shape and pose information simultaneously. Previously, template-free methods can reconstruct mammals [34, 35, 36, 37, 38] or birds [39] with arbitrary topology. Although these methods demonstrate high flexibility, their surface accuracy still has room for improvement. In this paper, we focus on template-based methods instead of template-free ones. Reconstructing shape templates has been studied in various types of animal families such as birds [28, 2], rodents [40], non-human primates [41] and quadrupeds [42]. For quadrupeds, the well-known SMAL [1] is built upon 41 scans of animal toys. Due to the limited geometry accuracy of SMAL for representing specific species, most previous methods stride over predicting SMAL parameters only and further enhance the geometry of horses and dogs. For example,  [4] designs the horse-specific hSMAL model and applies it to the video lameness detection problem. VAREN [3] further improves hSMAL with high quality horse scans. Similarly, the use of SMAL for dog mesh recovery is addressed by adding bone lengths control [11], by adding per-vertex deformation [10], by introducing a breed loss [9], by ground contact constraints [8], or by temporal avatar optimization [7, 43]. Though high quality reconstruction has been achieved on horses and dogs, end-to-end SMAL estimation has not been fully addressed on other challenging species. Animal3D [18] provides the first large-scale 3D benchmark for general quadruped SMAL estimation, yet neglects in-depth research on network design and training strategy.
@@
Transformer Based Human Mesh Recovery. The Transformer architecture [14] has revolutionized the field of Natural Language Processing (NLP) by enabling unprecedented accuracy and efficiency in a wide range of tasks. Inspired by its success in NLP, one core part of Transformer, i.e. self-attention, has been widely employed in human mesh recovery [44, 45, 46, 47].  [48] proposes ViT, which divides an image into patches as the input to Transformer. ViT has achieved state-of-the-art performance on several computer vision tasks including human mesh recovery using SMPL [44, 45, 46, 15, 47]. Among all these previous works, HMR2.0 [15] is a milestone which demonstrates the effectiveness of simply using a ViT backbone and large-scale datasets to achieve accurate mesh recovery and in-the-wild generalizability. Inspired by HMR2.0, HaMeR [16] employs a ViT backbone to achieve highly accurate hand mesh recovery, which is further extended to interacting hands [49]. Similarly, SMPLer-X [17] scales up expressive human pose and shape estimation using a ViT backbone and a combination of 32 datasets. Although impressive results have been achieved in human mesh recovery, the effect of employing a ViT backbone for animal pose and shape estimation remains unexplored.
@@
Synthetic Animal Training. Compared to human pose estimation which benefits from large-scale datasets, acquiring annotated images of animals is significantly more difficult. Synthetic datasets shall alleviate this issue by rendering the input and output simultaneously. Most previous methods only attempt to render RGB images [50, 51, 52, 8, 53, 18, 54] or depth images [55] using traditional CG pipelines, ignoring the image hallucination ability of generative AI models such as stable diffusion [56] or ControlNet [22]. Although traditional CG-based rendering achieves success in assisting network training, large-scale and high-quality CG assets are expensive to obtain, and the usability of rendered images is hindered by sophisticated lighting and shadow control. In contrast,  [57] uses 3D visual prompts and large language model (LLM) text prompts to generate diverse images from a static CAD model, pioneering a new direction. We extend their idea to animal image synthesis from dynamic SMAL animations to construct the CtrlAni3D dataset. [58] performs style-transfer from synthetic domain to real domain using ControlNet [22]. However, they rely on textured CAD assets instead of untextured SMAL, making it difficult to assist SMAL pose and shape estimation. Concurrent work GenZoo [59] generates realistic images with corresponding SMAL+ [60] parameters utilizing ControlNet in a similar way as ours, but they are specific to quadrupeds.
@@ III-A Preliminaries
The SMAL Model. The SMAL model [1], denoted as $\mathcal{M}(\beta,\theta,\gamma)$, is a 3D parametric shape model designed for quadrupedal mammals. The shape parameter $\beta\ \in\mathbb{R}^{41}$ is derived from 41 3D scans of various animal figurines, such as cats, dogs, horses, cows, and hippos. The pose parameter $\theta\in\mathbb{R}^{35\times 3}$ represents the rotation of each joint relative to its parent joint, expressed in terms of axis-angle. Controlled by $\beta$ and $\theta$, the SMAL model outputs a mesh with vertices $V\in\mathbb{R}^{3889\times 3}$ and faces $F\in\mathbb{N}^{7774\times 3}$ through a linear blend skinning (LBS) process. The animal body joints are regressed from vertices by $J\in\mathbb{R}^{35\times 3}=W\cdot V$, where $W\in\mathbb{R}^{35\times 3889}$ represents a linear mapping matrix.
@@
Camera Projection. Following HMR2.0 [15], $\pi(\cdot)$ represents the projection process of a weak-perspective camera model, determined by a translation vector $T\in\mathbb{R}^{3}$, a fixed focal length and thereby a fixed intrinsic matrix $K$. The global orientation $R$ is the rotation of SMAL’s root joint, therefore we ignore it here. Consequently, a 3D point $X$ is projected as a 2D point $x$ by $x=\pi(X)=\Pi(K(X+T))$, where $\Pi$ converts homogeneous coordinates $(u,v,d)^{T}$ to pixel coordinates $(u/d,v/d)^{T}$.

### v2

@@ I Introduction
Accurate estimation of animal pose and shape from images is essential for capturing animal behavior, biomechanics, and interactions with environment. It provides vital insights for diverse fields including animal welfare, agriculture, ecology, and life sciences. Integrating geometric and appearance information from widely varying species, such as birds and mammals, into a unified deep neural network represents a crucial step towards comprehensively interpreting the intricate and dynamic animal world. Significant efforts in this domain involve estimating pose and shape parameters using articulated animal templates. For example, the well-known SMAL model [zuffi20173d] primarily targets mammals, while the AVES parametric model [wang2021birds] focuses on birds. Although extensive research has been conducted on specific species or families, such as horses [zuffi2024varen, li2021hsmal, li2024dessie], zebras [zuffi2019three], dogs [Sabathier2024AnimalAR, ruegg2023bite, rueegg2022barc, li2021coarse, biggs2020wldo, biggs2019creatures] and birds [wang2021birds], investigations into other mammal species, including cats, cows or hippos, remain relatively underexplored, and few previous methods try to predict the meshes for animals of different classes (e.g., mammalia and aves) in a single network.
@@
Reconstructing multiple species within a single network is hindered by two primary challenges: the limited capacity of backbones and the scarcity of multi-species datasets annotated with corresponding parametric model labels. Recent advancements in human mesh recovery via the SMPL model [SMPL:2015] demonstrate that combining high-capacity Transformer [vaswani2017attention] backbones with large-scale datasets significantly enhances pose and shape estimation accuracy across diverse settings [goel2023humans, pavlakos2024reconstructing, cai2024smpler]. However, this effective paradigm has yet to be fully explored within animal studies. While  [xu2023animal3d] introduces the large-scale Animal3D dataset featuring SMAL annotations, their work primarily utilizes CNN-based networks like HMR [kanazawa2018end] and WLDO [biggs2020wldo].
@@
Figure 1: Statistics and representative samples of the CtrlAVES3D and CtrlAni3D datasets. (a) and (b) represent samples from the CtrlAVES3D dataset and the CtrlAni3D dataset, respectively. For each image pair, the left side displays the generated animal image whose background comes from either COCO  [lin2014microsoft] or AI-synthesis, and the right side presents the rendered mesh label. Note that the synthesis process naturally considers generating truncated images. (c) shows the statistics of both datasets. More details about these two datasets can be found in Table I, Table II and Sec. IV.
@@
To recap, our previous work, AniMer [lyu2025animer], proposes a systematic approach that pursues accurate animal pose and shape estimation by leveraging family-aware Transformer. The success of AniMer is built on two key scaling factors: a scaled backbone and a scaled dataset. Architecturally, AniMer integrates a high-capacity Vision Transformer (ViT) backbone with a Transformer-based decoder to regress SMAL parameters, a paradigm proven effective in human mesh recovery but previously untested for animal shape and pose estimation. To unify the discriminative understanding of diverse animal shapes within a single framework, AniMer introduces a novel animal family supervised contrastive learning scheme. It employs a learnable class token to explicitly enhance the model’s ability to distinguish between the unique shapes of different mammal families.
@@
A high-capacity model’s effectiveness is intrinsically linked to the quality and quantity of its training data. However, acquiring a large number of animal images with full 3D annotations is difficult, in which situations synthetic datasets can mitigate the scarcity of real data. The coarse texture quality and sophisticated lighting and shadow control hinder the usability of traditional Computer Graphics (CG) generated animal images. These limitations motivate our usage of ControlNet [zhang2023adding] to create high-quality AI hallucinated images conditioned on SMAL structures. Compared to CG-based, this principle effectively bridges the domain gap between parametric labels and images, and helps to create a novel, large-scale dataset named CtrlAni3D with minimal human labor involved. Specifically, we prompt ControlNet with textual descriptions of animal behaviors and rendered SMAL mask and depth images, resulting in highly realistic visual outputs, as illustrated in Fig. 1. To ensure dataset quality, we use SAM2.0 [ravi2024sam2] together with manual verification to filter inappropriate images. Finally, CtrlAni3D comprises 9711 images annotated with pixel-aligned SMAL meshes, and its scalable nature allows for further expansion. We end up with scaled training datasets by aggregating most available open-sourced quadrupedal datasets and our CtrlAni3D dataset, resulting in a comprehensive set of 41.3k images annotated with either 3D meshes or 2D keypoints.
@@
Despite these advances, the SMAL model is insufficient for animals like birds, whose skeletal structures fundamentally differ from mammals’. Training a shared model naively on both mammals and birds risks “averaging out” crucial anatomical features and suffers from severe depth ambiguities in single-view reconstruction, exacerbated by the fact that existing avian datasets provide only 2D annotations. Inspired by the success of Mixture-of-Experts (MoE) architectures [shazeer2017outrageously, jiang2024mixtral, xu2023vitpose++] in handling diverse tasks, we propose AniMer+, an extension of AniMer incorporating an MoE design to better unify different animal taxa. We adapt the ViT encoder by partitioning its fully-connected (FC) layers into taxa-specific layers (capturing expert quadruped/avian knowledge) and a taxa-shared layer (learning common semantic features). The taxa-specific layers process features from the corresponding taxa, while the taxa-shared layer receives features from all taxa. This architecture enables effective handling of both taxonomic groups within a single, unified framework. To enable robust 3D training for birds, we apply our generation pipeline to create CtrlAVES3D, the first large-scale avian dataset with 3D AVES model annotations, thereby mitigating the single-view ambiguity issue.
@@
To rigorously evaluate the efficacy of our AniMer and AniMer+ models, as well as our dataset generation pipeline, we conduct extensive empirical studies. For a clarification purpose, we name AniMer trained on mammals only, birds only and both mammals and birds respectively as AniMer-M, AniMer-A and AniMer-AM. Our findings indicate that AniMer, when trained on the complete multi-species mammal datasets (AniMer-M), significantly outperforms CNN-based methods such as HMR and WLDO [biggs2020wldo] which serve as baselines in the Animal3D research. Similarly, AniMer-A largely surpasses previous methods on bird datasets. Meanwhile, through ablation studies, we verify that the synthesized datasets enhance the generalizability of AniMer on the out-of-domain (OOD) Animal Kingdom dataset [Ng_2022_CVPR] and the OOD Cow Bird dataset [badger20203d], which are unseen during training. In addition, ablation studies also prove that the animal family-supervised contrastive learning scheme improves pose and shape estimation precision. Finally, the joint training of AniMer+ on both mammals and birds outperforms AniMer-AM, demonstrating the effectiveness of the MoE design.
@@
An earlier version of this work is AniMer [lyu2025animer], which was published in CVPR 2025. In this paper, we have significantly enhanced AniMer in three aspects. First, we create the first large-scale bird dataset with 3D annotations, named CtrlAVES3D, which fulfills the lack of 3D annotations in existing avian datasets and mitigates depth ambiguity in the monocular bird reconstruction task. Second, we propose AniMer+ by incorporating an MoE design to better unify different animal taxa, which outperforms AniMer when trained simultaneously on mammals and birds. Third, we conduct comparative and ablation experiments on avian species, establishing state-of-the-art performance across multiple avian datasets. Based on these extensions and in conjunction with the dataset generation pipeline, AniMer+ has the potential to generalize to any animal groups (e.g., fishes, insects, rodents or primates) that can be represented by parametric models.
@@ II Related Works
Pose and Shape Estimation for Mammals and Birds. Compared to monocular 2D/3D animal keypoint estimation methods [mathis2018deeplabcut, bala2020automated, pereira2022sleap, ye2024superanimal, dabhi20243d], monocular surface estimation methods can provide both shape and pose information simultaneously. Previously, template-free methods can reconstruct mammals [yang2022banmo, yao2022lassie, jakab2024farm3d, li2024learning, kaye2025dualpm] or birds [kanazawa2018learning] with arbitrary topology. Although these methods demonstrate high flexibility, their surface accuracy still has room for improvement. In this paper, we focus on template-based methods instead of template-free ones. Reconstructing shape templates has been studied in various types of animal families such as birds [badger20203d, wang2021birds], rodents [an2023three], non-human primates [neverova2020continuous] and quadrupeds [zuffi2018lions]. For quadrupeds, the well-known SMAL [zuffi20173d] is built upon 41 scans of animal toys. Due to the limited geometry accuracy of SMAL for representing specific species, most previous methods stride over predicting SMAL parameters only and further enhance the geometry of horses and dogs. For example,  [li2021hsmal] designs the horse-specific hSMAL model and applies it to the video lameness detection problem. VAREN [zuffi2024varen] further improves hSMAL with high quality horse scans. Similarly, the use of SMAL for dog mesh recovery is addressed by adding bone lengths control [biggs2020wldo], by adding per-vertex deformation [li2021coarse], by introducing a breed loss [rueegg2022barc], by ground contact constraints [ruegg2023bite], or by temporal avatar optimization [Sabathier2024AnimalAR, zhong20254d]. Though high quality reconstruction has been achieved on horses and dogs, end-to-end SMAL estimation has not been fully addressed on other challenging species. Animal3D [xu2023animal3d] provides the first large-scale 3D benchmark for general quadruped SMAL estimation, yet neglects in-depth research on network design and training strategy.
@@
Transformer Based Human Mesh Recovery. The Transformer architecture [vaswani2017attention] has revolutionized the field of Natural Language Processing (NLP) by enabling unprecedented accuracy and efficiency in a wide range of tasks. Inspired by its success in NLP, one core part of Transformer, i.e. self-attention, has been widely employed in human mesh recovery [kocabas2021pare, wan2021encoder, shen2023global, shin2024wham].  [dosovitskiy2020image] proposes ViT, which divides an image into patches as the input to Transformer. ViT has achieved state-of-the-art performance on several computer vision tasks including human mesh recovery using SMPL [kocabas2021pare, wan2021encoder, shen2023global, goel2023humans, shin2024wham]. Among all these previous works, HMR2.0 [goel2023humans] is a milestone which demonstrates the effectiveness of simply using a ViT backbone and large-scale datasets to achieve accurate mesh recovery and in-the-wild generalizability. Inspired by HMR2.0, HaMeR [pavlakos2024reconstructing] employs a ViT backbone to achieve highly accurate hand mesh recovery, which is further extended to interacting hands [lin20244dhands]. Similarly, SMPLer-X [cai2024smpler] scales up expressive human pose and shape estimation using a ViT backbone and a combination of 32 datasets. Although impressive results have been achieved in human mesh recovery, the effect of employing a ViT backbone for animal pose and shape estimation remains unexplored.
@@
Synthetic Animal Training. Compared to human pose estimation which benefits from large-scale datasets, acquiring annotated images of animals is significantly more difficult. Synthetic datasets shall alleviate this issue by rendering the input and output simultaneously. Most previous methods only attempt to render RGB images [cao2019cross, mu2020learning, plum2023replicant, ruegg2023bite, shooter2024digidogs, xu2023animal3d, zhong2025moremouse] or depth images [Kearney_2020_CVPR] using traditional CG pipelines, ignoring the image hallucination ability of generative AI models such as stable diffusion [rombach2022high] or ControlNet [zhang2023adding]. Although traditional CG-based rendering achieves success in assisting network training, large-scale and high-quality CG assets are expensive to obtain, and the usability of rendered images is hindered by sophisticated lighting and shadow control. In contrast,  [ma2023generating] uses 3D visual prompts and large language model (LLM) text prompts to generate diverse images from a static CAD model, pioneering a new direction. We extend their idea to animal image synthesis from dynamic SMAL animations to construct the CtrlAni3D dataset. [jiang2023spac] performs style-transfer from synthetic domain to real domain using ControlNet [zhang2023adding]. However, they rely on textured CAD assets instead of untextured SMAL, making it difficult to assist SMAL pose and shape estimation. Concurrent work GenZoo [niewiadomski2024generative] generates realistic images with corresponding SMAL+ [zuffi_eccv2024_awol] parameters utilizing ControlNet in a similar way as ours, but they are specific to quadrupeds.
@@ III-A Preliminaries
The SMAL Model. The SMAL model [zuffi20173d], denoted as $\mathcal{M}(\beta,\theta,\gamma)$, is a 3D parametric shape model designed for quadrupedal mammals. The shape parameter $\beta\ \in\mathbb{R}^{41}$ is derived from 41 3D scans of various animal figurines, such as cats, dogs, horses, cows, and hippos. The pose parameter $\theta\in\mathbb{R}^{35\times 3}$ represents the rotation of each joint relative to its parent joint, expressed in terms of axis-angle. Controlled by $\beta$ and $\theta$, the SMAL model outputs a mesh with vertices $V\in\mathbb{R}^{3889\times 3}$ and faces $F\in\mathbb{N}^{7774\times 3}$ through a linear blend skinning (LBS) process. The animal body joints are regressed from vertices by $J\in\mathbb{R}^{35\times 3}=W\cdot V$, where $W\in\mathbb{R}^{35\times 3889}$ represents a linear mapping matrix.
@@
Camera Projection. Following HMR2.0 [goel2023humans], $\pi(\cdot)$ represents the projection process of a weak-perspective camera model, determined by a translation vector $T\in\mathbb{R}^{3}$, a fixed focal length and thereby a fixed intrinsic matrix $K$. The global orientation $R$ is the rotation of SMAL’s root joint, therefore we ignore it here. Consequently, a 3D point $X$ is projected as a 2D point $x$ by $x=\pi(X)=\Pi(K(X+T))$, where $\Pi$ converts homogeneous coordinates $(u,v,d)^{T}$ to pixel coordinates $(u/d,v/d)^{T}$.